OpenStudy (anonymous):

solve???? x+y=5 x+2z=6 y+3z=4

OpenStudy (mathstudent55):

1. Solve the first equation for x. 2. Then substitute that into x of the second equation. 3. Now the second equation has only y and z. 4. The third equation also has only y and z. 5. Solve these two equations for y and z. 6. Then substitute y into the first equation to get x.

OpenStudy (anonymous):

how do you solve x+y=5 and get the answer for X?

OpenStudy (mathstudent55):

Subtract y from both sides.

OpenStudy (mathstudent55):

x + y = 5 - y - y ------------- x = 5 - y Now that you know that x = 5 - y, replace the x of the second equation with 5 - y.

OpenStudy (anonymous):

so igot z=5 y=-5 and x=10 is that right?

OpenStudy (mathstudent55):

That solution does not work in the second equation.

OpenStudy (anonymous):

use elimination process: x + y = 5 x + 2z = 6 ----------- (-) y - 2z = -1 y + 3z = 4 ----------- (-) -5z = -5 z = -5/(-5) = 1 y + 3z = 4 y + 3(1) = 4 y = 4 - 3 = 1 x + y = 5 x + 1 = 5 x = 5 - 1 = 4 so, x = 4, y = 1, z = 1

OpenStudy (anonymous):

Oh. Man I wonder what I did wrong???